OpenStudy (sethmc1828):

100= (1.75/1.015)+(1.75/1.0165^2)+(101.75/(1+x)^3) I keep getting .0169, but x is supposed to be .0175

OpenStudy (mathstudent55):

\(100= \dfrac{1.75}{1.015}+\dfrac{1.75}{1.0165^2}+\dfrac{101.75}{(1+x)^3} \) \(100= 3.4177864 +\dfrac{101.75}{(1+x)^3} \) \(96.582214 = \dfrac{101.75}{(1+x)^3} \) \(0.9492109 = \dfrac{1}{(1+x)^3} \) \(1.0535066 = (1 + x)^3\) \(1.01752266 = 1 + x\) \(0.01752266 = x\)